OpenStudy (anonymous):

If the slope is 3, then what is y in (5, y) (2, 2) ? myininaya!

OpenStudy (unklerhaukus):

11

OpenStudy (mertsj):

\[3=\frac{y-2}{5-2}\]

OpenStudy (mertsj):

Can you solve that?

OpenStudy (anonymous):

Yeah, I got to that part, but after that, I don't know how to solve it

OpenStudy (mertsj):

Well after you find y, you are done because the problem asks what is y if the slope is 3.

OpenStudy (anonymous):

so how would you find it?

OpenStudy (mertsj):

\[3=\frac{y-2}{3}\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (mertsj):

I subtracted 5-2 to get the 3 below y-2

OpenStudy (mertsj):

Now you should multiply both sides by 3

OpenStudy (anonymous):

OH sorry haha. okay

OpenStudy (mertsj):

\[3\times3=\frac{y-2}{3}\times3\]

OpenStudy (mertsj):

\[9=y-2\]

OpenStudy (mertsj):

\[11=y\]

OpenStudy (mertsj):

Are we good?